OpenStudy (mendicant_bias):

Now setting up another problem regarding calculating the displacement current of a parallel-plate capacitor, not sure if I'll need help yet, will be posting attempted workings beneath.

OpenStudy (mendicant_bias):

"A parallel-plate capacitor has square plates 10 cm on a side and 0.50 cm apart. The voltage across the plates is increasing at 220 V/ms. What is the displacement current across the capacitor?"

OpenStudy (mendicant_bias):

Alright, so what I ultimately need to be able to find is the Electric Flux. I'm guessing I can use the information given about the capacitor to calculate its electric field strength and then calculate its electric flux from that, and then go from flux to displacement current.

OpenStudy (mendicant_bias):

Alright, I found a formula for the electric field strength between two idealized, large parallel, charged plates. \[E=\frac{\sigma}{\epsilon}\]Sigma is charge density, and epsilon is permittivity. Found another equation expressing the voltage across the two plates to be equal to the electric field times the distance in between them, \[V = Ed.\]Since I have both the voltage and the distance between the two plates, I can solve for electric field strength: \[E = \frac{V}{d}=\frac{}{}\] Wait, no, the voltage isn't constant...

OpenStudy (mendicant_bias):

@blurbendy or @hba , I'm not sure if you're any good with Physics stuff, but I'm linking you in case you can help me; any help would be appreciated if you can.

OpenStudy (mendicant_bias):

Displacement current is always equal to the current that is charging the capacitor, apparently.

OpenStudy (mendicant_bias):

@ganeshie8 , are you any good with Electricity & Magnetism?

OpenStudy (mendicant_bias):

According to the Wiki page for Capacitors, http://upload.wikimedia.org/math/7/5/6/756d1aa6542ef0266a7aed62154c1c76.png

ganeshie8 (ganeshie8):

@dan815

OpenStudy (mendicant_bias):

\[I(t)=\frac{dQ(t)}{dt}=C\frac{dV(t)}{dt}\]

OpenStudy (mendicant_bias):

Alright, so it looks like we're given the rate of change of voltage with respect to time, so I'm guessing that's \[\frac{dV(t)}{dt}=\frac{220 V}{1 \ ms}\] (Speaking of, hey Ganeshie, do you know how I put text upright in the editor? I know there's a command like \textup in normal editors, but I can never get it to work on OS.)

OpenStudy (mendicant_bias):

\[C = \frac{\varepsilon A}{D}=\frac{k \varepsilon_{0} A}{D}\]

OpenStudy (mendicant_bias):

(But yeah, Ganeshie, do you know how I can put text upright in the editor?)

ganeshie8 (ganeshie8):

use `\rm` ?

OpenStudy (mendicant_bias):

\[C = \frac{(8.854 \cdot 10^{-12}F/ \rm m)(1)}{[(0.1 \ \rm m ) \cdot (0.1 \ \rm m)]}=\]

OpenStudy (mendicant_bias):

Thanks! I've been wanting to know how to do that for a long time, heh.

ganeshie8 (ganeshie8):

that doesn't look right

ganeshie8 (ganeshie8):

A is in numerator and d is in denominator

OpenStudy (mendicant_bias):

What doesn't look right about the calculation for C? epsilon naught and k are given values, the permittivity of free space, together, divided by-OH, yeah

OpenStudy (mendicant_bias):

Yep, you're right, denominator is distance, not area, thank you.

ganeshie8 (ganeshie8):

k=1 for air i think

OpenStudy (mendicant_bias):

So then it's just, \[\frac{8.854\cdot10^{-12}}{0.1\ \rm m}=8.854\cdot10^{-11}\]

ganeshie8 (ganeshie8):

`A parallel-plate capacitor has square plates 10 cm on a side and 0.50 cm apart.` A = 0.1^2 d = 0.005

OpenStudy (mendicant_bias):

UGUUUH thank you, I keep getting ahead of myself (and actually excited I appeared to have figured this out on my own) and am making mistakes everywhere...)

ganeshie8 (ganeshie8):

Nice :) I don't seem to remember much of these anymore :O

OpenStudy (mendicant_bias):

Yeah, alright, so trying to do this without making mistakes, heh, here we go:

OpenStudy (mendicant_bias):

Due to the fact that for a capacitor,\[I_{D}=I,\]or displacement current is equivalent to the current running *through* a capacitor, \[I_{D}=I=C \frac{dV(t)}{dt}=\frac{\varepsilon A}{D}\frac{dV(t)}{dt}\]

OpenStudy (mendicant_bias):

\[\frac{\varepsilon A}{D} \frac{dV(t)}{dt}=\frac{8.854\cdot10^{-12}F/ \rm m}{0.5\cdot10^{-2}\ \rm m}\bigg(220\cdot 10^{-3} \rm V/\rm ms\Bigg)\]

OpenStudy (mendicant_bias):

Whoops, the units in the dV/dt should be just volts per second due to the times ten to the negative three.

OpenStudy (mendicant_bias):

\[=3.89576\cdot10^{-10}\] is what I got, my algebra might be off. I hope not.

ganeshie8 (ganeshie8):

\[\frac{\varepsilon A}{D} \frac{dV(t)}{dt}=\frac{(8.854\cdot10^{-12}F/ \rm m )(0.1 )^2}{0.5\cdot10^{-2}\ \rm m}\bigg(220~ \rm V/\rm ms\Bigg) \]

ganeshie8 (ganeshie8):

3.896 nA

OpenStudy (mendicant_bias):

Whoah, lol, I just....totally forgot a term in there. Thank you!

ganeshie8 (ganeshie8):

np:) the currentlooks very small hmm

OpenStudy (mendicant_bias):

(Uh oh?)